# Nonstandard Universes

### v1

@@ 5.1 Definitions and Basic Properties
To verify that a transfer map $*:\mathbb{U}\to\mathbb{V}$ for a set $X$ is a $\kappa$-enlargement, it suffices to consider sets $\mathcal{A}\in\mathbb{U}$ in (a) of of Theorem 5.2.
To verify that a transfer map $*:\mathbb{U}\to\mathbb{V}$ for a set $X$ is $\kappa$-saturated, it suffices to consider sets $\mathcal{A}$ which are subsets of standard sets — and thus in $\mathbb{V}$ — in (b) of of Theorem 5.2.
Proof: Suppose that * satisfies (i) of Definition 5.1 for sets $\mathcal{A}\in\mathbb{U}$. Let $\mathcal{A}^{\prime}\subseteq\mathbb{U}$ be a set of cardinality $<\kappa$ which satisfies the f.i.p. Choose $A_{0}\in\mathcal{A}^{\prime}$, and let $\mathcal{A}:=\{A\cap A_{0}:A\in\mathcal{A}^{\prime}\}$. Observe that $|\mathcal{A}|<\kappa$, that $\mathcal{A}$ satisfies the f.i.p. as well, and that since $\mathcal{A}\subseteq\mathcal{P}(A_{0})$ we have $\mathcal{A}\in\mathbb{U}$. By assumption, $\bigcap{}^{\sigma}\mathcal{A}\neq\varnothing$. Now as ${}^{*}(A\cap A_{0})\subseteq{}^{*}A$ for all $A\in\mathcal{A}^{\prime}$, we have $\varnothing\neq\bigcap{}^{\sigma}\mathcal{A}=\bigcap{}^{\sigma}\mathcal{A}^{\prime}$.
Next, suppose that * satisfies (ii) of Definition 5.1 for $\mathcal{B}\subseteq{}^{*}\mathcal{A}$, where $\mathcal{A}\in\mathbb{U}$. Let $\mathcal{B}^{\prime}\subseteq{}^{*}\mathbb{U}$ be a family of internal sets of cardinality $<\kappa$ which satisfies the f.i.p. Choose $B_{0}\in\mathcal{B}^{\prime}$, and let :=$\mathcal{B}:=\{B\cap B_{0}:B\in\mathcal{B}^{\prime}\}$. Observe that $|\mathcal{B}|<\kappa$, that $\mathcal{B}$ has the f.i.p. also, and that $\mathcal{B}\subseteq{}^{*}\mathbb{U}\cap\mathcal{P}(B_{0})$. By Lemma 3.8, $P:={}^{*}\mathbb{U}\cap\mathcal{P}(B_{0})$ is internal, and hence there is some transitive $\mathcal{A}\in\mathbb{U}$ such that $P\in{}^{*}\mathcal{A}$. Then also $\mathcal{B}\subseteq{}^{*}\mathcal{A}$, and hence by assumption we have that $\bigcap\mathcal{B}\neq 0$. Hence $\varnothing\neq\bigcap\mathcal{B}=\bigcap\mathcal{B}^{\prime}$.
@@
It follows from Theorem 5.10 that $*:\mathbb{U}\to\mathbb{V}$ is an enlargement if and only if it is a $|\mathbb{U}|$-enlargement, as $\mathcal{A}\in\mathbb{U}$ implies $|\mathcal{A}|<|\mathbb{U}|$.
@@
Suppose that $B\in{}^{*}\mathbb{U}$ is an infinite internal set. Then for $b\in B$, the set $B_{b}:=\{c\in B:c\neq b\}$ $=B-\{b\}$ is internal, by the internal definition principle and the transitivity of ${}^{*}\mathbb{U}$. Clearly :=$\mathcal{B}:=\{B_{b}:b\in B\}$ has the f.i.p., yet $\bigcap\mathcal{B}=\varnothing$. Hence * cannot be a $|B|^{+}$-saturated extension. In particular, if $*:\mathbb{U}\to\mathbb{V}$ is a $\kappa$-saturated transfer map for an infinite set $X$ of atoms, then necessarily $|{}^{*}X|\geq\kappa$. It is therefore impossible to find an extension which is $\kappa$-saturated for every cardinal $\kappa$.
@@
If $*:\mathbb{U}\to\mathbb{V}$ is $\kappa$-saturated and $B\in\mathbb{V}$ has cardinality $|B|<\kappa$, then $B$ is external.
@@
If $*:\mathbb{U}\to\mathbb{V}$ is $\kappa$-saturated, it is a $\kappa$-enlargement. Hence a polysaturated extension is an enlargement.
Proof: If $\mathcal{A}\subseteq\mathcal{U}$ is a family of sets with the f.i.p., then $\mathcal{B}:={}^{\sigma}\mathcal{A}$ is a family of internal (indeed, standard) sets with the f.i.p. For, given $A_{1},\dots,A_{n}\in\mathcal{A}$, transfer of the bounded sentence $\exists x\in A_{1}(x\in A_{2}\land\dots\land x\in A_{n})$ shows that ${}^{*}A_{1}\cap\dots\cap{}^{*}A_{n}\neq\varnothing$. Clearly $|\mathcal{B}|=|\mathcal{A}|$.
@@ 5.2 Enlargements, Concurrency and Hyperfinite Approximation
A transfer map $*:\mathbb{U}\to\mathbb{V}$ for an infinite set $X$ is $\omega_{1}$-enlargement if and only if it is a nonstandard embedding.
Proof: ()($\Rightarrow$)( ⇒ ): Suppose that * is an $\omega_{1}$-enlargement. To show that * is a nonstandard embedding, it suffices to show that there is a countable $C\in\mathbb{U}$ such that ${}^{\sigma}C\subsetneqq{}^{*}C$. So let $C\subseteq X$ be countable. Since $\mathcal{P}(X)\in\mathbb{U}$ and $\mathbb{U}$ is transitive, we have $C\in\mathbb{U}$. Let $\mathcal{C}:=\{C-\{c\}:c\in C\}$. Then $\mathcal{C}$ has the f.i.p. and $|\mathcal{C}|<\omega_{1}$, so $\bigcap{}^{\sigma}\mathcal{C}\neq\varnothing$. Pick $c_{0}\in\bigcap{}^{\sigma}\mathcal{C}$. Then $c_{0}\in{}^{*}(C-\{c\})={}^{*}C-\{{}^{*}c\}$ for every $c\in C$. Hence $c_{0}\in{}^{*}C-{}^{\sigma}C$.
()($\Leftarrow$)( ⇐ ): Now assume that * is a nonstandard embedding. Let $\mathcal{A}\subseteq\mathbb{U}$ be a family of sets with the f.i.p. such that $|\mathcal{A}|<\omega_{1}$, i.e. $|\mathcal{A}|$ is countable. Let $\mathcal{A}=\{A_{1},A_{2},\dots\}$ be an enumeration of $\mathcal{A}$. Without loss of generality (by renaming elements if necessary), we may assume that $\mathbb{N}\subseteq X$. Let $f:\mathbb{N}\twoheadrightarrow\mathcal{P}(A_{1}):n\mapsto\bigcap_{k\leq n}A_{k}$. Then $f\in\mathbb{U}$, and $\varnothing\not\in\text{ran}f$. By transfer, $\varnothing\not\in\text{ran}\,{}^{*}f$, since ${}^{*}\varnothing=\varnothing$ and ${}^{*}(\text{ran}f)=\text{ran}\,{}^{*}f$. In particular, for all $n\in{}^{*}\mathbb{N}$, ${}^{*}f(n)\neq\varnothing$. Now fix an arbitrary $n\in\mathbb{N}$, and note that
@@
(Recall ${}^{*}x=x$ for all $x\in X$, and thus ${}^{*}n=n$.) Since * is a nonstandard embedding, it follows from Theorem 3.16 that there exists $m_{0}\in{}^{*}\mathbb{N}-{}^{\sigma}\mathbb{N}$ for which we have both ${}^{*}f(m_{0})\neq\varnothing$ and $m_{0}\geq n$. It follows that ${}^{*}f(m_{0})\subseteq{}^{*}A_{n}$. Since $n\in\mathbb{N}$ was arbitrary, we have that $\varnothing\neq f(m_{0})\subseteq\bigcap_{n\in\mathbb{N}}{}^{*}A_{n}=\bigcap{}^{\sigma}\mathcal{A}$. This shows that * is an $\omega_{1}$-enlargement.
@@
For a set $A$, and important example of a concurrent relation is the set
$R:=\{(a,F):a\in F\in\mathcal{P}^{<\omega}(A)\}.$
@@
Suppose that $*:\mathbb{U}\to\mathbb{V}$ is a transfer map for a set $X$. Then the following are equivalent:
@@
    $*:\mathbb{U}\to\mathbb{V}$ is a $\kappa$-enlargement.
@@
(ii)$\Rightarrow$ (iii): Given a set $A\in\mathbb{U}$ with $|A|<\kappa$, define the concurrent relation $R$ by
@@
(iii) $\Rightarrow$ (i): Let $\mathcal{A}\in\mathbb{U}$ have the f.i.p. with $|\mathcal{A}|<\kappa$, and let $T\in\mathbb{U}$ be transitive so that $\mathcal{A}\subseteq T$. Then $A\subseteq T$ for all $A\in\mathcal{A}$, and hence
@@ 5.3 Saturation and Concurrency
Suppose that $*:\mathbb{U}\to\mathbb{V}$ is a nonstandard transfer map. Then the following are equivalent:
@@
    $*:\mathbb{U}\to\mathbb{V}$ is $\kappa$-saturated.
@@
(ii) $\Rightarrow$ (i): Suppose that $\mathcal{A}$ is a family of internal sets with the f.i.p., where $|\mathcal{A}|<\kappa$. We must show that $\bigcap\mathcal{A}\neq\varnothing$. Fix $A_{0}\in\mathcal{A}$. By replacing each $A\in\mathcal{A}$ by $A\cap A_{0}$, we may assume that $A\subseteq A_{0}$ for each $A\in\mathcal{A}$. Thus $\mathcal{A}\subseteq\mathcal{P}(A_{0})\cap{}^{*}\mathbb{U}$. By Lemma 3.10, $P(A_{0})\cap{}^{*}\mathbb{U}$ is an internal set. Consider the relation
$R:=\{(A,a)\in(P(A_{0})\cap{}^{*}\mathbb{U})\times A_{0}:a\in A\}.$
@@ 5.4 Comprehensiveness
Let $\kappa$ be an infinite cardinal. A transfer map $*:\mathbb{U}\to\mathbb{V}$ is said to be $\kappa$-comprehensive if and only if for any sets $A,B\in\mathbb{U}$ such that $|A|<\kappa$, and any map $f:A\to{}^{*}B$, there is an internal function ${}^{+}\!f:{}^{*}A\to{}^{*}B$ with the property that ${}^{+}\!f({}^{*}a)=f(a)$ for all $a\in A$.
$*:\mathbb{U}\to\mathbb{V}$ is said to be comprehensive if and only if it is $\kappa$-comprehensive for every cardinal $\kappa$, or equivalently, if it is $|\mathbb{U}|^{+}$–comprehensive.
$*:\mathbb{U}\to\mathbb{V}$ is said to be countably comprehensive if and only if it is $\omega_{1}$-comprehensive.
@@
If a transfer map $*:\mathbb{U}\to\mathbb{V}$ is $\kappa$-comprehensive, then for any set $A\in\mathbb{U}$ of cardinality $|A|<\kappa$, and any internal set $B\in\mathbb{V}$, if $f:A\to B$, then there is an internal function ${}^{+}\!f:{}^{*}A\to B$ with the property that ${}^{+}\!f({}^{*}a)=f(a)$ for all $a\in A$.
@@
Suppose that $*:\mathbb{U}\to\mathbb{V}$ is a countably comprehensive transfer map, where $\mathbb{U}$ is a universe over a set $X$ that contains $\mathbb{N}$. Suppose also that $B$ is internal and that $(b_{n}:n\in\mathbb{N})$ is a sequence of members of $B$. Then there is an internal sequence $(b_{n}:n\in{}^{*}\mathbb{N})$ which extends the sequence $(b_{n}:n\in\mathbb{N})$.
@@
Suppose that $*:\mathbb{U}\to\mathbb{V}$ is a $\kappa$-saturated nonstandard embedding. Then * is $\kappa$-comprehensive.
@@
Let :=$\mathcal{B}:=\{B_{A}:a\in A\}$, so that $|\mathcal{B}|<\kappa$. To apply $\kappa$-saturation, we must show that $\mathcal{B}$ has the f.i.p. Now given for each $n\in\mathbb{N}$, we have that
@@
[…]
@@
Suppose that $*:\mathbb{U}\to\mathbb{V}$ is a nonstandard embedding, where $\mathbb{U}$ is a universe over a set $X$ that contains $\mathbb{N}$. The following are equivalent:
@@
Proof: This follows directly from Theorem 5.15.

### v2

@@ 5.1 Definitions and Basic Properties
To verify that a transfer map $\mathbb{U}\to\mathbb{V}$ for a set $X$ is a $\kappa$-enlargement, it suffices to consider sets $\mathcal{A}\in\mathbb{U}$ in (a) of of Theorem 5.2.
To verify that a transfer map $\mathbb{U}\to\mathbb{V}$ for a set $X$ is $\kappa$-saturated, it suffices to consider sets $\mathcal{A}$ which are subsets of standard sets — and thus in $\mathbb{V}$ — in (b) of of Theorem 5.2.
Proof: Suppose that * satisfies (i) of Definition 5.1 for sets $\mathcal{A}\in\mathbb{U}$. Let $\mathcal{A}^{\prime}\subseteq\mathbb{U}$ be a set of cardinality $\kappa$ which satisfies the f.i.p. Choose $A_{0}\in\mathcal{A}^{\prime}$, and let $\mathcal{A}:=\{A\cap A_{0}:A\in\mathcal{A}^{\prime}\}$. Observe that $|\mathcal{A}|<\kappa$, that $\mathcal{A}$ satisfies the f.i.p. as well, and that since $\mathcal{A}\subseteq\mathcal{P}(A_{0})$ we have $\mathcal{A}\in\mathbb{U}$. By assumption, $\bigcap{}^{\sigma}\mathcal{A}\neq\varnothing$. Now as $^{*}(A\cap A_{0})\subseteq{}^{*}A$ for all $A\in\mathcal{A}^{\prime}$, we have $\varnothing\neq\bigcap{}^{\sigma}\mathcal{A}=\bigcap{}^{\sigma}\mathcal{A}^{\prime}$.
Next, suppose that * satisfies (ii) of Definition 5.1 for $\mathcal{B}\subseteq{}^{*}\mathcal{A}$, where $\mathcal{A}\in\mathbb{U}$. Let $\mathcal{B}^{\prime}\subseteq{}^{*}\mathbb{U}$ be a family of internal sets of cardinality $\kappa$ which satisfies the f.i.p. Choose $B_{0}\in\mathcal{B}^{\prime}$, and let $\mathcal{B}:=\{B\cap B_{0}:B\in\mathcal{B}^{\prime}\}$. Observe that $|\mathcal{B}|<\kappa$, that $\mathcal{B}$ has the f.i.p. also, and that $\mathcal{B}\subseteq{}^{*}\mathbb{U}\cap\mathcal{P}(B_{0})$. By Lemma 3.9, $P:={}^{*}\mathbb{U}\cap\mathcal{P}(B_{0})$ is internal, and hence there is some transitive $\mathcal{A}\in\mathbb{U}$ such that $P\in{}^{*}\mathcal{A}$. Then also $\mathcal{B}\subseteq{}^{*}\mathcal{A}$, and hence by assumption we have that $\bigcap\mathcal{B}\neq 0$. Hence $\varnothing\neq\bigcap\mathcal{B}=\bigcap\mathcal{B}$$^{\prime}$.
@@
It follows from Theorem 5.10 that $\mathbb{U}\to\mathbb{V}$ is an enlargement if and only if it is a $|\mathbb{U}|$-enlargement, as $\mathcal{A}\in\mathbb{U}$ implies $|\mathcal{A}|<|\mathbb{U}|$.
@@
Suppose that $B\in{}^{*}\mathbb{U}$ is an infinite internal set. Then for $b\in B$, the set $B_{b}:=\{c\in B:c\neq b\}$ $B-\{b\}$ is internal, by the internal definition principle and the transitivity of $^{*}\mathbb{U}$. Clearly $\mathcal{B}:=\{B_{b}:b\in B\}$ has the f.i.p., yet $\bigcap\mathcal{B}=\varnothing$. Hence * cannot be a $|B|^{+}$-saturated extension. In particular, if $\mathbb{U}\to\mathbb{V}$ is a $\kappa$-saturated transfer map for an infinite set $X$ of atoms, then necessarily $|{}^{*}X|\geq\kappa$. It is therefore impossible to find an extension which is $\kappa$-saturated for every cardinal $\kappa$.
@@
If $\mathbb{U}\to\mathbb{V}$ is $\kappa$-saturated and $B\in\mathbb{V}$ has cardinality $|B|<\kappa$, then $B$ is external.
@@
If $\mathbb{U}\to\mathbb{V}$ is $\kappa$-saturated, it is a $\kappa$-enlargement. Hence a polysaturated extension is an enlargement.
Proof: If $\mathcal{A}\subseteq\mathbb{U}$ is a family of sets with the f.i.p., then $\mathcal{B}:={}^{\sigma}\mathcal{A}$ is a family of internal (indeed, standard) sets with the f.i.p. For, given $A_{1},\dots,A_{n}\in\mathcal{A}$, transfer of the bounded sentence $\exists x\in A_{1}(x\in A_{2}\land\dots\land x\in A_{n})$ shows that $^{*}A_{1}\cap\dots\cap{}^{*}A_{n}\neq\varnothing$. Clearly $|\mathcal{B}|=|\mathcal{A}|$.
@@ 5.2 Enlargements, Concurrency and Hyperfinite Approximation
A transfer map $\mathbb{U}\to\mathbb{V}$ for an infinite set $X$ is $\omega_{1}$-enlargement if and only if it is a nonstandard embedding.
Proof: $(\Rightarrow)$: Suppose that * is an $\omega_{1}$-enlargement. To show that * is a nonstandard embedding, it suffices to show that there is a countable $C\in\mathbb{U}$ such that $^{\sigma}C\subsetneqq{}^{*}C$. So let $C\subseteq X$ be countable. Since $\mathcal{P}(X)\in\mathbb{U}$ and $\mathbb{U}$ is transitive, we have $C\in\mathbb{U}$. Let $\mathcal{C}:=\{C-\{c\}:c\in C\}$. Then $\mathcal{C}$ has the f.i.p. and $|\mathcal{C}|<\omega_{1}$, so $\bigcap{}^{\sigma}\mathcal{C}\neq\varnothing$. Pick $c_{0}\in\bigcap{}^{\sigma}\mathcal{C}$. Then $c_{0}\in{}^{*}(C-\{c\})={}^{*}C-\{{}^{*}c\}$ for every $c\in C$. Hence $c_{0}\in{}^{*}C-{}^{\sigma}C$.
$(\Leftarrow)$: Now assume that * is a nonstandard embedding. Let $\mathcal{A}\subseteq\mathbb{U}$ be a family of sets with the f.i.p. such that $|\mathcal{A}|<\omega_{1}$, i.e. $|\mathcal{A}|$ is countable. Let $\mathcal{A}=\{A_{1},A_{2},\dots\}$ be an enumeration of $\mathcal{A}$. Without loss of generality (by renaming elements if necessary), we may assume that $\mathbb{N}\subseteq X$. Let $f:\mathbb{N}\twoheadrightarrow\mathcal{P}(A_{1}):n\mapsto\bigcap_{k\leq n}A_{k}$. Then $f\in\mathbb{U}$, and $\varnothing\not\in\text{ran}f$. By transfer, $\varnothing\not\in\text{ran}\,{}^{*}f$, since $^{*}\varnothing=\varnothing$ and $^{*}(\text{ran}f)=\text{ran}\,{}^{*}f$. In particular, for all $n\in{}^{*}\mathbb{N}$, $^{*}f(n)\neq\varnothing$. Now fix an arbitrary $n\in\mathbb{N}$, and note that
@@
(Recall $^{*}x=x$ for all $x\in X$, and thus $^{*}n=n$.) Since * is a nonstandard embedding, it follows from Theorem 3.16 that there exists $m_{0}\in{}^{*}\mathbb{N}-{}^{\sigma}\mathbb{N}$, i.e. an infinite $m_{0}$. Then we have both $^{*}f(m_{0})\neq\varnothing$ and $m_{0}\geq n$. It follows that $^{*}f(m_{0})\subseteq{}^{*}A_{n}$. Since $n\in\mathbb{N}$ was arbitrary, we have that $\varnothing\neq f(m_{0})\subseteq\bigcap_{n\in\mathbb{N}}{}^{*}A_{n}=\bigcap{}^{\sigma}\mathcal{A}$. This shows that * is an $\omega_{1}$-enlargement.
@@
For a set $A$, an important example of a concurrent relation is the set
$R:=\{(a,F):a\in F\in\mathcal{P}^{<\omega}(A)\}\subseteq A\times\mathcal{P}^{<\omega}(A).$
@@
Suppose that $\mathbb{U}\to\mathbb{V}$ is a transfer map for a set $X$. Then the following are equivalent:
@@
    $\mathbb{U}\to\mathbb{V}$ is a $\kappa$-enlargement.
@@
[…]
@@
(iii) $\Rightarrow$ (i): We use Theorem 5.3. Let $\mathcal{A}\in\mathbb{U}$ have the f.i.p. with $|\mathcal{A}|<\kappa$, and let $T\in\mathbb{U}$ be transitive so that $\mathcal{A}\subseteq T$. Then $A\subseteq T$ for all $A\in\mathcal{A}$, and hence
@@ 5.3 Saturation and Concurrency
Suppose that $\mathbb{U}\to\mathbb{V}$ is a nonstandard embedding. Then the following are equivalent:
@@
    $\mathbb{U}\to\mathbb{V}$ is $\kappa$-saturated.
@@
(ii) $\Rightarrow$ (i): Suppose that $\mathcal{A}$ is a family of internal sets with the f.i.p., where $|\mathcal{A}|<\kappa$. We must show that $\bigcap\mathcal{A}\neq\varnothing$. Fix $A_{0}\in\mathcal{A}$. By replacing each $A\in\mathcal{A}$ by $A\cap A_{0}$, we may assume that $A\subseteq A_{0}$ for each $A\in\mathcal{A}$. Thus $\mathcal{A}\subseteq\mathcal{P}(A_{0})\cap{}^{*}\mathbb{U}$. By Lemma 3.9, $\mathcal{P}(A_{0})\cap{}^{*}\mathbb{U}$ is an internal set. Consider the relation
$R:=\{(A,a)\in(\mathcal{P}(A_{0})\cap{}^{*}\mathbb{U})\times A_{0}:a\in A\}.$
@@ 5.4 Comprehensiveness
Let $\kappa$ be an infinite cardinal. A transfer map $\mathbb{U}\to\mathbb{V}$ is said to be $\kappa$-comprehensive if and only if for any sets $A,B\in\mathbb{U}$ such that $|A|<\kappa$, and any map $f:A\to{}^{*}B$, there is an internal function $^{+}\!f:{}^{*}A\to{}^{*}B$ with the property that $^{+}\!f({}^{*}a)=f(a)$ for all $a\in A$.
$\mathbb{U}\to\mathbb{V}$ is said to be comprehensive if and only if it is $\kappa$-comprehensive for every cardinal $\kappa$, or equivalently, if it is $|\mathbb{U}|^{+}$–comprehensive.
$\mathbb{U}\to\mathbb{V}$ is said to be countably comprehensive if and only if it is $\omega_{1}$-comprehensive.
@@
If a transfer map $\mathbb{U}\to\mathbb{V}$ is $\kappa$-comprehensive, then for any set $A\in\mathbb{U}$ of cardinality $|A|<\kappa$, and any internal set $B\in\mathbb{V}$, if $f:A\to B$, then there is an internal function $^{+}\!f:{}^{*}A\to B$ with the property that $^{+}\!f({}^{*}a)=f(a)$ for all $a\in A$.
@@
Suppose that $\mathbb{U}\to\mathbb{V}$ is a countably comprehensive transfer map, where $\mathbb{U}$ is a universe over a set $X$ that contains $\mathbb{N}$. Suppose also that $B$ is internal and that $(b_{n}:n\in\mathbb{N})$ is a sequence of members of $B$. Then there is an internal sequence $(b_{n}:n\in{}^{*}\mathbb{N})$ which extends the sequence $(b_{n}:n\in\mathbb{N})$.
@@
Suppose that $\mathbb{U}\to\mathbb{V}$ is a $\kappa$-saturated nonstandard embedding. Then * is $\kappa$-comprehensive.
@@
Let $\mathcal{B}:=\{B_{a}:a\in A\}$, so that $|\mathcal{B}|<\kappa$. To apply $\kappa$-saturation, we must show that $\mathcal{B}$ has the f.i.p. Now for each $n\in\mathbb{N}$, we have that
@@
$\mathbb{V}\vDash\text{$g$ is a function}\land\text{\rm dom}(g)={}^{*}A\land({}^{*}a_{1},f(a_{1}))\in g\land\dots\land({}^{*}a_{n},f(a_{n}))\in g,$
@@
Suppose that $\mathbb{U}\to\mathbb{V}$ is a nonstandard embedding, where $\mathbb{U}$ is a universe over a set $X$ that contains $\mathbb{N}$. The following are equivalent:
@@
Proof: (i) $\Rightarrow$ (ii): This follows directly from Theorem 5.15.